### v3

# Disentangling CLIP for Multi-Object Perception

Abstract
Vision-language models like CLIP excel at recognizing the single, prominent object in a scene. However, they struggle in complex scenes containing multiple objects. We identify a fundamental reason behind this limitation: VLMs features space exhibits significant semantic entanglement, where features of one class contain substantial information about other unrelated classes, a phenomenon we term mutual feature information (MFI). This entanglement becomes evident during class-specific queries, as unrelated objects are activated alongside the queried class. To address this limitation, we propose DCLIP, a framework that disentangles CLIP features using two complementary objectives: a novel MFI Loss that orthogonalizes the text (class) features to reduce inter-class similarity, and the Asymmetric Loss (ASL) that aligns image features with the disentangled text features. Our experiment demonstrates that DCLIP reduces inter-class feature similarity by 30% compared to CLIP, leading to significant performance gains on multi-label recognition (MLR) and zero-shot semantic segmentation (ZS3). In MLR, DCLIP outperforms SOTA approaches on VOC2007 and COCO-14 while using 75% fewer parameters, and surpasses SOTA ZS3 methods by 3.4 mIoU on VOC2012 and 2.8 mIoU on COCO-17. These results establish feature disentanglement as a critical factor for effective multi-object perception in vision-language models.
1 Introduction
Vision-language models (VLMs) such as CLIP have emerged as powerful tools for understanding visual content through natural language supervision. CLIP trains on a massive dataset of 400 million image-text pairs and demonstrates impressive performance in recognizing the salient object in the image, retrieving similar images from large datasets, and answering image-related natural language queries. However, an important question arises: do these impressive capabilities transfer when CLIP processes images containing multiple objects? As illustrated in Fig. 1, CLIP often struggles in such scenarios, failing to accurately recognize and localize all the objects present in the image. In this work, we investigate and identify the causes of this limitation and propose a framework that enables VLMs to handle complex images with multiple objects.
Figure 1: Illustration of Feature Entanglement in CLIP. Given the query “Human”, CLIP exhibits feature entanglement, spuriously activating regions corresponding to other objects (dogs and horse). In contrast, DCLIP (ours) accurately focuses only on the human. The rightmost plot quantifies this improvement through reduced cosine similarity between class features. Takeaway: By significantly decreasing inter-class similarity (Human-Horse: 0.84 $\rightarrow$ 0.42 [50% $\downarrow$], Human-Dog: 0.80 $\rightarrow$ 0.28 [65% $\downarrow$], DCLIP yields more disentangled features, crucial for precise multi-object perception.
In our investigation, we systematically analyze CLIP’s features and identify two key factors contributing to this limitation. First, the spatial pooling operation in the visual encoder’s final layer, while sufficient for identifying the prominent object, eliminates crucial spatial information needed to recognize and locate multiple distinct objects in an image. Second, and more importantly, we discover significant entanglement between class features in the joint vision-language space, a phenomenon we term mutual feature information (MFI). This entanglement becomes apparent during class-specific queries: regions containing unrelated objects activate alongside the target class. For eg, as illustrated in Fig. 1, “dog” and “horse” regions also activate when we query “human.” This activation pattern strongly correlates with the high similarity scores between class features (0.84 for human-horse and 0.80 for human-dog). We extend this analysis to the classes in VOC [11] and COCO [19] datasets, where we observe average feature similarities of 0.77 in VOC and 0.69 in COCO (Tab. 5), confirming substantial feature entanglement in CLIP’s feature space.
To address the feature entanglement in VLMs, we propose DCLIP, a framework that disentangles class features using two complementary objectives. We draw inspiration from the redundancy reduction principle [2], which states that sensory systems (eg. brain) recode input information such that redundancy is minimized without losing useful information, and extend it to the vision-language domain to reduce MFI. While previous approaches have focused on architectural modifications [30, 18, 3] or prompt engineering [27, 23] to adapt VLMs for multi-object settings, they do not address the fundamental feature entanglement problem. DCLIP targets this root cause through our novel MFI Loss that explicitly disentangles text features by orthogonalizing them to reduce inter-class similarity, while the Asymmetric Loss (ASL) ensures proper cross-modal alignment by projecting image features to align with the disentangled text features. This joint training using MFI loss (separates text features) and ASL loss (aligns text and image features) results in a semantically disentangled feature space that significantly improves multi-object perception capabilities.
We evaluate DCLIP on multi-label recognition (MLR) and zero-shot semantic segmentation (ZS3). using established benchmarks: COCO-14 and VOC2007 [11] for MLR, and VOC2012 [11] and COCO-17 [19] for ZS3. Our experimental results demonstrate that DCLIP reduces inter-class feature similarity by an average of 30% compared to CLIP across these datasets, leading to favorable MLR performance over SOTA methods on VOC2007 and the challenging COCO-14 dataset, while requiring 75% fewer parameters. For ZS3, DCLIP surpasses SOTA VLM methods, demonstrating that reducing mutual feature information (MFI) is crucial for multi-object perception.
The main contributions of this work are:
  • We identify a critical limitation of VLMs for multi-object perception: mutual information between class features (MFI) negatively impacts recognition and localization performance
  • We propose DCLIP, a framework that disentangles class features through our novel MFI and the ASL loss, creating a disentangled feature space while preserving image-text alignment
  • The disentangled feature space is suitable for both MLR and ZS3, and outperforms SOTA on both tasks while requiring 75% fewer training parameters (trained only for MLR)
2 Related Work
(a) VOC Dataset
(b) COCO Dataset
Figure 2: Inter-Class Feature Similarity Analysis. Comparison of class feature similarities between CLIP and DCLIP (our) across the VOC and COCO datasets. The heatmaps visualize the inter-class cosine similarity, where darker blue indicates higher similarity. Takeaway: DCLIP features demonstrate better separation as seen by reduced inter-class (off-diagonal entries) similarity values.
Vision-Language Models for Multi-Object Perception. Vision-language models (VLMs) trained with contrastive losses [22, 16] are challenging to adapt for multi-object settings for two reasons: (1) their reliance on global feature aggregation, which ignores local information. (2) Using the softmax operation in their training loss biases them toward single-object settings.
Recognition. Early efforts to adapt VLMs for recognition centered on learning prompts as classifiers for visual features [31]. These methods were extended to multi-label settings by learning multiple prompts for each class [27, 15, 23]. Subsequent works incorporated co-occurrence information to make predictions interdependent [9, 24]. In contrast, our approach does not rely on prompt learning or co-occurrence modeling during pre-training. Furthermore, our features are adaptable to tasks beyond multi-label recognition.
Localization. Early approaches addressed localization by training image segmentation models and using VLMs to label the segmented regions [17]. Later methods introduced pre-training setups that combined vision-language alignment with mask distillation to enhance localization [10]. Recent works adapted features for localization without additional training by leveraging the spatial properties preserved in the value projection of CLIP’s transformer-style aggregation [30]. CLIP Surgery [18] identified consistent noisy activations across classes and reduced them by subtracting average features from class-specific features [18], though the cause of these activations remains unclear. GEM generalized this concept to vision transformers [3]. We use the finding that value projection preserves spatial information. We further improve value projection by disentangling class features.
Recoding information. Shannon proposed that optimal information transmission involves designing codes with minimum entropy [26]. The redundancy reduction principle extended this idea to neuroscience, suggesting that sensory systems recode information to reduce redundancy with minimal loss [2]. This principle has since been applied to many recent works, including image compression [1] and more popularly in representation learning [21, 5, 29, 14, 13, 6]. While our loss function shares structural similarities with representation learning methods (a similarity and contrastive term), our method differs as follows: (1) DCLIP uniquely refines pre-trained VLM features, directly manipulating an existing, semantically rich space rather than learning representations from scratch (2) Crucially, our MFI loss operates on the fixed set of class text embeddings to directly reduce inter-class semantic similarity. This contrasts fundamentally with instance-discriminative contrastive losses that rely on intra-sample invariance (3) DCLIP is adapted for multi-object perception in VLMs, a setting prior works do not address.
Figure 3: DCLIP Overview. Given image and class names in the dataset, CLIP extracts image and text features, which are projected by respective projectors to a disentangled space while preserving local image information. To reduce mutual feature information (MFI) between class features, we propose MFI loss that enforces the self-similarity matrix of projected text features to approximate an identity matrix, effectively reducing inter-class feature dependencies (Sec. 3.3). To align the image and the separated text features we the ASL loss in a multi-label recognition setup (Sec. 3.4). This setup aggregates the projected image and text features to obtain predicted logits (Sec. 3.4). The predicted logits are trained with ground truth labels using the widely used asymmetric loss (ASL) [25]. Our training loss combines the ASL and MFI loss, the only trainable components are the projectors. During inference, we freeze CLIP’s image and text encoders, along with our projectors for multi-label recognition and zero-shot semantic segmentation tasks.
3 DCLIP
3.1 Analysis of Feature Entanglement in CLIP
Before we introduce DCLIP, we first analyze the feature entanglement phenomenon in CLIP that limits its multi-object perception capabilities. Our examination of pairwise cosine similarity between class text features reveals unexpectedly high inter-class similarities in CLIP’s feature space, with average values of 0.77 for VOC and 0.69 for COCO classes (Fig. 2). Strikingly, even semantically distinct categories like “human” and “horse” show similarity scores as high as 0.84, far exceeding what their semantic relationship would suggest. This feature entanglement manifests visually when querying for specific classes in complex scenes. As shown in Fig. 1, when querying for “human,” CLIP erroneously activates “dog” and “horse” regions, a direct consequence of their entangled feature representations. The distribution of cosine similarity in Fig. 4 confirms this is not an isolated issue but rather a pervasive problem affecting the majority of class pairs. We attribute this entanglement to CLIP’s contrastive training objective, which optimizes for global image-text alignment without explicitly constraining separation between different class features. Additionally, the global pooling operation mixes features across the entire image, which further exacerbates feature entanglement. Our analysis establishes a clear correlation between feature entanglement reduction and improved performance on multi-object perception (Fig. 6). As mutual feature information (MFI) decreases, both multi-label recognition and semantic segmentation performance consistently improve, confirming that addressing feature entanglement is crucial for enhancing CLIP’s multi-object perception. This insight motivates our DCLIP framework, which maintains CLIP’s rich semantic knowledge while explicitly reducing mutual information between class features to improve multi-object perception.
DCLIP Overview. Based on our analysis, we propose DCLIP, a framework that disentangles class features to enable effective multi-object perception. DCLIP leverages a pre-trained CLIP model ($f_{\theta}$), which comprises an image encoder ($f_{\theta,\text{img}}$) and a text encoder ($f_{\theta,\text{text}}$), both parameterized by $\theta$. All CLIP parameters ($\theta$) are frozen for all experiments. The DCLIP framework operates on multi-label dataset $\mathcal{D}=\{\mathbf{x}_{i}\,\mathbf{y}_{i}\}_{i=1}^{|\mathcal{D}|}$, where each image $\mathbf{x}_{i}$ is associated with a label vector $\mathbf{y}_{i}\in\{0,1\}^{N}$ indicating the presence of objects from multiple classes within our label space consisting of $N$ distinct classes $\{C_{j}\}_{j=1}^{N}$. DCLIP consists of three components: (1) Feature extraction and Projection, where we extract CLIP local features and project them into a disentangled space (Sec. 3.2), (2) Defining novel MFI Loss for disentangling text features (Sec. 3.3), and (3) Using ASL to align image features to the disentangled text features (Sec. 3.4).
3.2 Feature Extraction and Projection
We use CLIP as our feature extractor. Its image encoder ($f_{\theta,\text{img}}$) performs spatial pooling in the final layer, aggregating features from local regions into a $d$-dimensional vector for the input image $x_{i}$. However, as elaborated in Sec. 3.1, this pooling step, suppresses the contribution of less prominent objects, making it unsuitable for images with multiple objects. To mitigate this, we remove the final pooling layer of $f_{\theta,\text{img}}$ to preserve class-specific information across local regions. Then the encoder output for input ($\mathbf{x}_{i}$) is ${f}_{\theta,img}(\mathbf{x_{i}})=\mathbf{z}_{i}\in\mathbb{R}^{H\times W\times d}$ , where $H$ and $W$ are the spatial dimensions. The text encoder remains unchanged. We use a fixed pair of positive and negative ($\mathbf{txt_{j,+}}$, $\mathbf{txt_{j,-}}$) prompts for each class $j$ as input to the text encoder. $\mathbf{txt_{j,+}}$ indicates the presence of the class in the image, while the $\mathbf{txt_{j,-}}$ indicates its absence. Passing these prompts through the text encoder produces ${f}_{\theta,text}(\mathbf{txt_{i}})=\mathbf{t}_{i}$.
Following the feature extraction process, the resulting features (image ($\mathbf{z}_{i}$), text ($\mathbf{t}_{i}$)) lie in CLIP’s feature space. As previously discussed in Sec. 3.1, this space exhibits significant feature entanglement and is not suitable for multi-object perception. To mitigate this, we project $\mathbf{z}_{i}$ and $\mathbf{t}_{i}$ into a new disentangled space using learnable projectors ($h_{\phi}:h_{\phi,\text{img}}$ and $h_{\phi,\text{text}}$), parameterized by weights $\phi$. These projectors map $\mathbf{z}_{i}$ and $\mathbf{t}_{i}$ from original space ($d$-dim) to a new disentangled space ($d^{\prime}$-dim). Specifically, $h_{\phi,\text{img}}$ transforms $\mathbf{z}_{i}\to\mathbf{z^{\prime}}_{i}$ ($\mathbb{R}^{H\times W\times d}\to\mathbb{R}^{H\times W\times d^{\prime}}$) while preserving the spatial dimensions ($H,W$). In addition, $h_{\phi,\text{text}}$ maps $\mathbf{t}_{i}\to\mathbf{t^{\prime}}_{i}$ ($\mathbb{R}^{d}\to\mathbb{R}^{d^{\prime}}$).
3.3 MFI Loss
We design our projected feature space to reduce mutual feature information (MFI) between class features. To reduce MFI, we require isolated inputs for each class, as MFI measures the information shared between features of these isolated inputs. For images with multiple objects, obtaining isolated class inputs is challenging. For eg, an image containing both a horse and a dog would require computationally intensive segmentation methods to separate these objects. In contrast, text inputs are inherently isolated in our framework, as we already have individual class names (e.g., “horse” and “dog”) available in the dataset. We leverage this natural isolation of text inputs by extracting text features, projecting them into our shared space to get $\mathbf{t^{\prime}}$, and applying our proposed MFI loss to reduce information overlap between classes. Instead of using computationally expensive object segmentation for images, we propose an elegant solution that aligns these separated text features with image features using ASL loss (Sec. 3.4). Our MFI loss function is formulated as:
$\mathcal{L}_{\text{MFI}}=\underbrace{\sum_{i=1}\left(\mathbf{S}_{ii}-1\right)^{2}}_{\text{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}Collapse Prevention}}}+\lambda\underbrace{\sum_{i=1}\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}\mathbf{S}_{ij}^{2}}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}MFI Reduction}}}$ (1)
where $\mathbf{S}$ is the self-similarity matrix and is given by $\mathbf{S}_{ij}=\mathbf{t^{\prime}}_{i}^{\top}\mathbf{t^{\prime}}_{j}$ where $\mathbf{t^{\prime}}_{i},\mathbf{t^{\prime}}_{j}$ are the $i$-th and $j$-th column vectors of $\mathbf{t^{\prime}}$ (i.e., $\mathbf{t^{\prime}}_{i},\mathbf{t^{\prime}}_{j}\in\mathbb{R}^{1\times d^{\prime}}$). In this formulation, $\lambda$ is the hyperparameter that addresses the imbalance in the loss arising from the larger number of MFI reduction terms in $\mathbf{S}$ compared to the collapse prevention terms. The MFI loss minimizes the inter-class similarity $\mathbf{S}_{ij}$ ($i\neq j$) while simultaneously preserving high intra-class $\mathbf{S}_{ii}$ to prevent feature collapse. We provide the detailed connection of MFI to the Information Bottleneck principle in Appendix Sec. A.
(a) VOC Dataset
(b) COCO Dataset
Figure 4: Distribution of Inter-class Feature Similarities. Histograms of cosine similarity values between class features for CLIP (blue) and DCLIP (red) across VOC (left) and COCO (right) datasets. Takeaway: DCLIP consistently shifts the distribution toward lower similarity values on both datasets, demonstrating significant feature separation between classes.
3.4 Image-Text Alignment with ASL
MLR Formulation. MLR task involves identifying the subset of classes $\mathcal{C}_{i}\subseteq\{C_{1},C_{2},\ldots,C_{N}\}$ associated with the image $\mathbf{x}_{i}$. The goal is to learn a mapping function $g:\mathbf{x}_{i}\rightarrow\{-1,1\}^{N}$, that maps input images to $1$ if the class is present and $-1$ if the class is absent in the image.
We align $\mathbf{t^{\prime}_{i}}$ and $\mathbf{z^{\prime}_{i}}$ using the ASL loss in the MLR setup, eliminating the need for image segmentation. For each location $(h,w)$ in $(\mathbf{z^{\prime}}_{i})$, we detect the presence or absence of a class $j$, by computing the cosine similarity with positive text features ($\mathbf{t^{\prime}_{j,+}}$) and negative text features ($\mathbf{t^{\prime}_{j,-}}$). Higher similarity with the positive text features indicates the presence of the class, while a higher similarity with the negative text features indicates its absence. This process generates class-specific positive and negative similarity maps across the image. Following [27, 24, 23], we apply softmax to these maps to increase focus on local regions containing the class. See Appendix B for details. We then aggregate across the local regions to obtain the logits ${p_{i}}$ for the image. We train the setup with the widely used Asymmetric Loss function (ASL) [25], which addresses the significant imbalance between negative and positive examples in multi-label dataset. The ASL loss is given by:
$\mathcal{L}_{ASL}(p_{i}^{j})=\begin{cases}\left(1-p_{i}^{j}\right)^{\gamma_{+}}\log\left(p_{i}^{j}\right),&\text{if }y_{i}^{j}=1,\\ \left(p_{i,\delta}^{j}\right)^{\gamma_{-}}\log\left(1-p_{i,\delta}^{j}\right),&\text{otherwise }\end{cases}$ (2)
where $p_{i}^{j}$ represents the corresponding prediction associated with label $y_{i}^{j}$, $i$ represents the image and $j$ represents the class. $p_{i,\delta}^{j}=\max(\hat{y}-\delta,0)$, with $\delta$ is the shifting parameter defined in ASL.
Table 1: Comparison on multi-label recognition (MLR). We compare the performance (mAP) and training efficiency (parameters, GPU hours) of our approach with SOTA VLM-based MLR methods on VOC2007 and COCO-14 datasets. DCLIP performs favorably over SOTA on VOC2007, and on the challenging COCO dataset, it outperforms SOTA while requiring only one-fourth of the parameters. The reported mAP values represent the mean and standard deviation over 5 runs of DCLIP. red and blue indicate the best and the second best performance.
Methods VOC2007 COCO-14
Param($\downarrow$) GPU hrs($\downarrow$) mAP($\uparrow$) Param($\downarrow$) GPU hrs($\downarrow$) mAP($\uparrow$)
DualCoOp [27] 0.3M 3.6 94.2 1.3M 16 83.6
SCPNet [9] - $>$3.6 94.3 3.4M 26 84.4
TAI-DPT [12] $>$0.3M - - $>$1.3M - 84.5
DualCoOp++ [15] 0.4M - 94.9 1.5M - 85.1
MLR-GCN [24] 0.4M 3.6 94.4 1.3M - -
PositiveCoOp [23] 0.2M 3 94.4 0.8M 15 84.7
Ours 0.4M 3 95.2 $\pm$ 0.2 0.4M 13 85.4 $\pm$ 0.2
Training. Our training objective is composed of two components: (1) mutual feature information loss that enforces the separation between class text features and (2) Asymmetric loss function [25], designed for MLR that aligns the image features and text features to obtain predictions for an image.
$\mathcal{L}_{\text{DCLIP}}=\mathcal{L}_{\text{ASL}}+\alpha\mathcal{L}_{\text{MFI}}$ (3)
where $\alpha$ controls the relative importance of the two objectives.
4 Experiments
In this section, we describe the datasets, evaluation metrics, implementation details, and performance analysis for multi-label recognition (MLR) and zero-shot semantic segmentation (ZS3).
4.1 Datasets and Metrics
1) Pre-training with MLR: We evaluate the MLR performance using mean-Average Precision (mAP) on datasets: (1) COCO-14 [19] contains 80 classes with 82,081 training and 40,504 validation images. Following recent works [27, 24, 23], we train on the training set and evaluate on the validation set. (2) VOC2007 [11] is another widely used MLR dataset containing 20 classes with 9,963 images. Following [27, 24, 23], we use the train-val set for training and the test set for evaluation.
2) Zero-Shot Semantic Segmentation (ZS3): We use image and text projectors trained on the COCO-14 dataset and evaluate ZS3 using the mIoU metric on the following datasets: (1) PASCAL VOC 2012 [11] includes segmentation masks for the 20 classes in VOC2007. Following works [18, 3], we evaluate on the validation set. (2) COCO-2017 [19] includes segmentation masks for the 80 classes in COCO-14. Following [18, 3], we evaluate on the validation set.
4.2 Implementation Details
We use CLIP’s [22] original pre-trained encoder weights for all our experiments and keep them frozen. Consistent with popular MLR and ZS3 literature, we use a ResNet-based visual encoder and the standard transformer for text encoding [27, 9, 15, 23, 24, 12, 18, 20]. We conduct all experiments on a single RTX A4000 GPU.
During the pre-training stage with the MLR setup (Sec. 3.4), we follow the settings and hyperparameters from recent works [27, 24, 23]. This includes resizing images to $448$, applying Cutout [8] and RandAugment [7] for augmentation. Our projectors ($h_{\phi}$) are implemented as multi-layer perceptrons (MLPs). Specifically, the image projector follows a [512 $\rightarrow$ 256] architecture, while the text projector is designed as [512 $\rightarrow$ 384 $\rightarrow$ 256] with batch normalization and ReLU. We train both projectors with stochastic gradient descent (SGD) using an initial learning rate of 0.002, which is reduced by cosine annealing. We train the DCLIP setup (ASL + MFI loss) for 50 epochs with a batch size of 32. We follow [27, 24, 23], and use ASL hyperparameters in eq. 2 as $\gamma_{-}=2$, $\gamma_{+}=1$ and $\delta$ = 0.05. We set $\lambda$ = 0.2 and $\alpha$ = $7\mathrm{e}{-5}$ when pre-trained with COCO-14 in eq. 1.
For Zero-Shot Semantic Segmentation, we adopt the v-v attention described in [18] that prevents inversion of activation commonly observed in CLIP. We then add our pre-trained projectors to CLIP. To obtain the segmentation mask, we compute the cosine similarity between locally projected image features ($\mathbf{z^{\prime}}$) and projected text features for all classes in the dataset. We use the template “A photo of a {classname}.” Lastly, we use bilinear interpolation to upsample the mask to the input image size.
Table 2: Comparison on zero-shot semantic segmentation (ZS3). We compare DCLIP with other SOTA baselines across three semantic segmentation datasets using the mIoU metric. The abbreviations are as follows: Loc Ann. + FT: local annotations and fine-tuning, Bkgd: include background class, No Bkgd: ignore background class, MR: MFI Reduction(%), red and blue indicate the best and the second best performance.
Method Loc Ann. VOC12 COCO-17
+ FT Bkgd Bkgd No Bkgd
SPNet [28] 15.6 - -
ZS3Net [4] 17.7 - -
CLIP-ES [20] 75.0 - -
CLIP [22] 14.1 3.9 5.6
CLIPSurgery [18] 17.5 13.0 22.9
CLIP-VV [18] 32.6 19.9 35.5
Ours (MR = 24.9) 36.0 22.7 37.8
Table 3: Leveraging DCLIP Segments for Enhanced Zero-Shot Multi-Label Recognition. CLIP recognition capabilities often degrade in images with multiple objects. We demonstrate that incorporating object segments generated from DCLIP provides additional evidence to the global image features for zero-shot multi-label recognition (ZS-MLR) on VOC2007 and COCO-14. The resulting performance gains highlight DCLIP’s ability to extract single-object information, enabling robust zero-shot recognition with CLIP in multi-object images.
Dataset Backbone CLIP (mAP) Ours (mAP)
VOC2007 RN 101 78.73 80.71
RN 50 76.20 79.87
COCO-14 RN 101 50.10 52.00
RN 50 47.30 50.15
4.3 Results
Multi-Label Recognition. We primarily compare DCLIP with other SOTA VLM-based MLR approaches. In Tab. 1, we present a detailed comparison of the performance (mAP-averaged over five runs), number of training parameters, and GPU hours required by each method on the VOC2007 [11] and COCO-14 [19] datasets. For VOC 2007, we observe that DCLIP performs favorably over DualCoOp++[15], requiring equal parameters but fewer training hours. Additionally, on the more challenging COCO-14 dataset, DCLIP outperforms DualCoOp++ while requiring 75% fewer training parameters and fewer training hours on an NVIDIA A4000 GPU.
Zero-Shot Semantic Segmentation. We categorize our comparisons into two groups. The first group includes approaches that use local annotations (segmentation masks) to fine-tune the network [28, 4, 20]. The second group do not use any local annotations [22, 18]. Our approach belongs to the second group, as we do not use any form of local annotations Our results are summarized in Tab. 3. Following prior works [28, 4, 20, 18], we report mIoU values for VOC2012 by including the background as a class. Following [3], we use a threshold of 0.85 to identify the background. Our approach performs favorably over CLIP Surgery by 18.5 mIoU and CLIP-VV by 3.4 mIoU on VOC2012. On COCO-14, our method outperforms CLIP Surgery and CLIP-VV by 9.7 and 2.8 mIoU with background included, and by 14.9 and 2.3 mIoU without background, respectively.
5 Analysis
Image Oracle CLIP [22] CS [18] CLIP-VV [18] DCLIP (ours)
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 5: Qualitative ZS3 Comparison. Visualization of ZS3 results for CLIP [22], CLIP Surgery (CS) [18], CLIP-VV [18], and DCLIP (ours) across multiple categories. The red regions in Oracle show the queried classes. The heatmaps for different methods show activation regions for each queried class, where darker red indicates strongly activated regions. DCLIP (ours) produces more separated activations, leading to improved class localization.
Feature Disentanglement and its Impact. This section examines both quantitative and qualitative impacts of class feature disentanglement. In Fig. 2, we visualize self-similarity matrices for class text features, revealing that DCLIP achieves substantially lower inter-class similarity (off-diagonal values) than baseline CLIP. The distribution of similarity values in Fig. 4 further illustrates how DCLIP shifts feature representations toward reduced inter-class similarity. We quantitatively assess this disentanglement in Tab. 5 using average inter-class similarity, where DCLIP consistently demonstrates lower values across datasets, confirming effective feature separation. Crucially, Fig. 6 establishes a clear inverse relationship between MFI and performance: as MFI decreases, indicating greater feature disentanglement, we observe gains in both multi-label recognition and zero-shot semantic segmentation. Our ablation experiments provide additional validation, showing that removing the MFI Loss component for separation results in a significant 0.6 mAP reduction in MLR task performance (Tab. 5). These results strongly confirm that disentangling class features is both necessary and beneficial for improving performance for multi-object perception.
Segmentation-driven Zero-Shot Multi-Label Recognition (ZS-MLR). To evaluate whether DCLIP produces meaningful segments, we reformulate the multi-label recognition problem into a single-label recognition problem, a domain more suitable for CLIP. Specifically, we use two predictions: global and local. We pass the input image directly through CLIP to obtain its global predictions. However, as discussed in Sec. 1, these predictions are often dominated by more prominent objects in the image, ignoring smaller objects, which leads to poor zero-shot MLR performance. To address this limitation, we introduce local predictions. We segment the image into multiple regions (ideally corresponding to individual objects) using DCLIP. Each segment is then processed independently through CLIP, and the predictions from all segments are aggregated. Finally, we combine the global and local predictions to obtain the zero-shot scores for the image. We evaluate the ZS-MLR performance on the VOC2007 and COCO-14 datasets, with results presented in Tab. 3. The results demonstrate that DCLIP provides meaningful information (individual object segments) to improve CLIP’s zero-shot multi-label recognition capabilities.
Ablation. We study the sensitivity of MLR performance to the sensitivity of hyperparameters and projector choices. (1) $\boldsymbol{\alpha}$ (ASL–MFI Trade-off): Fig. 7 shows that varying $\alpha$, the coefficient that balances the ASL and MFI loss, has minimal impact on performance, indicating low sensitivity to $\alpha$. (2) $\boldsymbol{\lambda}$ (MFI Loss Weighting): We varied $\lambda$, which controls the trade-off between collapse prevention and redundancy reduction in the MFI loss (Eq. 1), across the range $[0.02,0.2]$. Performance remained stable throughout, suggesting insensitivity to $\lambda$ (Appendix Sec E). (3) Projector Dimension: Increasing the dimension of the projector beyond 256 dimensions led to no notable performance gains (Fig. 7).
Table 4: Quantitative MFI Reduction. MFI values are reported across VOC and COCO datasets. DCLIP significantly reduces MFI.
Method VOC COCO
CLIP 0.77 0.69
DCLIP(Ours) 0.50 0.52
$\Delta$ (%) 34.8 24.9
Table 5: MFI Loss Ablation. Without MFI Loss, we observe multi-label recognition performance drops by 0.6 mAP on the COCO-14 dataset.
Method ASL Loss MFI Loss mAP
Ours 84.8
85.4
Figure 6: Performance vs. MFI Reduction. Increase in class feature separation (i.e., MFI decreases) improves performance on multi-label recognition (mAP) and zero-shot semantic segmentation (mIoU) tasks.
Figure 7: Hyperparameter Ablations. 1) DCLIP demonstrates stability across $\alpha$ variations with fluctuations within 2mAP. 2) Performance improves with increased projector dimensionality, saturating beyond 256 dimensions.
6 Conclusions and Broader Impact
In conclusion, our work identifies that high mutual feature information (MFI) between class features impairs CLIP’s ability for multi-object perception. To address this, we propose DCLIP, a framework that disentangles CLIP features using our proposed MFI loss and the ASL loss. Experiments across multiple benchmarks show that reducing feature entanglement significantly improves multi-label recognition and zero-shot segmentation performance. These results establish feature disentanglement as essential for adapting VLMs to complex visual scenes without task-specific tuning. Limitations of our approach include handling fine-grained subcategories within the same superclass (e.g., different dog breeds or species of birds). This limitation partially stems from CLIP’s inherent limitation to fine-grained discrimination. Future work could explore adaptive disentanglement strategies that operate at multiple semantic levels to better support hierarchical visual concepts. Positive and Negative Societal Impacts: Our work enhances VLMs perception capabilities in multi-object scenes with applications in robotics and medical imaging. While we do not have direct negative societal impacts, perception technologies could be misused in surveillance systems or automated profiling applications. These concerns are not unique to our approach but apply broadly to computer vision systems, and appropriate governance frameworks should be developed alongside technical advances.